OpenStudy (anonymous):

What is the equation of the line with an x-intercept of -1 and a y-intercept of 2?

OpenStudy (anonymous):

use, \[(y-y1)=m(x-x1)\] m designates your slope, x1 ,y1 are your intercepts.

OpenStudy (anonymous):

So, (0,-1) (0,2)?

OpenStudy (fibonaccichick666):

The x intercept is -1, think about the point you want to use

OpenStudy (shamim):

(-1,0),(0,2)

OpenStudy (shamim):

so, x1=-1,y1=0,x2=0,y2=2

OpenStudy (shamim):

ur required equation of line is y-y1=((y1-y2)/(x1-x2))(x-x1)